Consider the following utility function of 2 goods, x and y: U(x,y)= - [(x-10)2 + (y-10)2]; x,y≥0

The prices of good x and y is 10 and 20 respectively. The income is denoted by m.

a. Draw the indifference curves for the utility function and use arrows to explain in which direction utility increases or decreases.

b. Find the consumption bundle that maximizes utility for the consumer.

c. Find the Engel curve.
Answer #1

(QJ UCXV) - - [ CX-10)2 + (Y-10)2 ] - # - [ 2x-20 + 24-20] - - 2x + go-24+Q0 u(x, 4) = -2x-2y + 40 = -2(x+4) 740 MRS = dolox. Utility utility increases decreases L x u= 0 ~ U=10 a up (=20 s w U= 28 PU= 32 97 U-361 / 1 2 3 4 5 6 7 8 9 10 X U=40 (b)But duc To - in utility function mini of (x,y) combination glucs, maximum utility and possible minimum consumption of x=0 and